OpenStudy (anonymous):

Can someone help with this question! Please! At STP, gases have a volume of 22.4 L per mole. RTP is room temperature (25 °C) and pressure (1 atm). What is the molar volume of a gas at RTP?

OpenStudy (wolfe8):

Can you use the formula PV=nRT?

OpenStudy (anonymous):

Yes, but when I do. I get 18.7L

OpenStudy (anonymous):

I don't know what I'm doing wrong?

OpenStudy (wolfe8):

Ok can you show me what you did? And make sure you're using the R that corresponds to the units you are using. There are at least 2 different R's

OpenStudy (anonymous):

okay one second

OpenStudy (anonymous):

Okay I re-did it and now I get 20.55mols/L so Idk PV=nRT n= RT/PV n=(8.31)(298K)/(101.31kpa)(22.4L) n=1.09mols 22.4L/1.09mol=20.55

OpenStudy (wolfe8):

Is it right? Logically yes since a lower temperature reduces pressure and volume.

OpenStudy (anonymous):

The answer is 24.5L

OpenStudy (anonymous):

but I have no idea how or what I'm doing wrong

OpenStudy (wolfe8):

Oooh wait. At STP the pressure is 1 atm. Since at RTP the pressure is the same, the only change here are for the temperature and volume. So we can use the law V1/T1=V2/T2

OpenStudy (anonymous):

Oh. Okay I'll try that. Thanks. :)

OpenStudy (anonymous):

it works! Thank you so much!

OpenStudy (wolfe8):

You're welcome. Sorry I didn't see that earlier. Would've made this shorter. Any other questions?